# Global calibration of large-scale photonic integrated circuits

### v1

@@ Abstract
The advancing maturity of photonic integrated circuit (PIC) fabrication technology enables the high integration of an increasing number of optical components onto a single chip. With the incremental circuit complexity, the calibration of active phase shifters in a large-scale PIC becomes a crucially important issue. The traditional one-by-one calibration techniques encounter significant hurdles with the propagation of calibration errors, and achieving the decoupling of all phase shifters for independent calibration is not straightforward. To address this issue, we propose a machine-learning approach for globally calibrating the large-scale PIC. Our method utilizes a custom network to simultaneously learn the nonlinear phase-current relations for all thermo-optic phase shifters on the PIC by minimizing the negative likelihood of the measurement datasets. Moreover, the reflectivities of all static beamsplitter components can also be synchronizedly extracted using this calibration method. As an example, a quantum walk PIC with a circuit depth of 12 is calibrated, and a programmable discrete-time quantum walk is experimentally demonstrated. These results will greatly benefit the applications of large-scale PICs in photonic quantum information processing.
@@ I Introduction
The traditional calibration method requires the decoupling of all phase shifters followed by individual calibration for each one [19, 20, 21, 22, 23]. However, this approach isn’t ideal for large-scale PICs due to two main challenges. Firstly, while the traditional method can be used to calibrate the standard triangular “Rech” PICs [24, 25], rectangular “Clements” PICs [26, 27, 28, 29] and PICs with cascaded phase shifters [30, 31] by following a specific decoupling order, it is difficult to find an effective and direct way to decouple all phase shifters in the intricate mesh structures. Secondly, deviations in the calibration values of phase shifters in preceding levels may also spread errors in the calibration of subsequent phase shifters. As a result, calibration errors accumulate progressively with the increasing number of phase shifters, eventually becoming significant and non-negligible.
To address these issues, we present an efficient calibration method developed to globally calibrate all optical components on the PICs, thereby mitigating the accumulation of errors inherent in the traditional one-by-one calibration approach. Our method is based on a white-box machine learning technique [32] that creates a virtual duplicate of the physical PICs by constructing a custom network, accurately modelling its real unitary processes. Using the gradient-descent approach, we can determine the optimal values of parameters that need to be calibrated, ensuring that the outputs of the network closely resemble the distributions of measurement data. As an example, on a large-scale quantum walk (QW) PIC with a circuit depth of 12, we find that the trained custom network enables the simultaneous extraction of not only the nonlinear phase-current relations of all the thermo-optic phase shifters but also the reflectivities of all the beamsplitters. Moreover, we further experimentally demonstrate the silicon PIC calibration and implement the on-chip QW [33, 34, 35, 36, 37, 38, 39, 40, 41].
This article is organized as follows: In Sec. II, we will introduce our custom-architecture machine-learning calibration method, provide a concise overview of the QW PIC, and evaluate the performance of our calibration method using synthetic datasets featuring Gaussian noise. In Sec. III, we will showcase the calibration results of a silicon PIC utilizing noisy experimental datasets and discuss the performance regarding the experimental implementation of QW. Finally, Sec. IV presents the summary of our work and engages in a discussion.
@@ II Theoretical Idea
In this section, we will detail our method for achieving precise and fast calibration of PICs. Based on the idea of white-box machine learning, we customize the hidden layers situated between the input and output layers, as depicted in Fig. 1. The entire custom network is trained using gradient-based optimization techniques to align its virtual process with the real physical process of the PICs, thereby allowing it to estimate the chip parameters that need to be calibrated.
II.1 Machine-learning-based calibration method
Figure 1: The conceptual scheme of a white-box machine learning architecture. The input data consists of the current distribution applied to the thermo-optic phase shifters. The first hidden layer of our custom network transforms the distribution of the current $\bm{I}$ into the distribution of the phase shift $\bm{\phi}$, with a subsequent layer outputting the parameterized final state |$\ket{\psi_{\text{f}}}$ for a given initial state |$\ket{\psi_{\text{i}}}$. The third hidden layer provides the corresponding measurement probability distribution of the final state $|\langle v|\psi_{\text{f}}\rangle|^{2}$ in the computational basis |$\ket{v}$. The fourth hidden layer introduces a parameter $\bm{\eta}$ into this probability distribution, which accounts for the relative fiber-chip coupling efficiency between output ports. The last layer renormalizes the probability distribution so that the sum of the probabilities over all output ports equals one, and a parameterized network distribution $\bm{P{{}_{\text{ML}}}}(\bm{\lambda})$ is generated thereafter.
@@ II.1 Machine-learning-based calibration method
where $\sin^{2}\theta$ denotes the reflectivity of the beamsplitter. The phase shifter provides a relative phase difference $\phi$ between the two output paths of the beamsplitter, and its matrix representation is $\hat{U}_{\text{PS}}=\text{diag}\{1,e^{i\phi}\}$. For the widely-used thermo-optic phase shifter, the relative phase $\phi$ can be tuned by the current $I$ applied to the phase shifter, according to the following formula [13]:
@@
where $a$ and $b$ are hardware parameters to be determined. Then, the unitary process $\hat{U}(\bm{a},\bm{b},\bm{\theta})$ of PICs is the matrix product of its individual components. Moreover, considering the mode-field diameter mismatch between the output ports of the PICs and the optical fiber arrays, it is necessary to introduce another parameter, $\bm{\eta}$, to indicate the fiber-chip coupling efficiency. Consequently, $\bm{\lambda}=\{\bm{a},\bm{b},\bm{\theta},\bm{\eta}\}$ are all undetermined parameters for a linear photonic circuit.
In contrast to the traditional one-by-one localized calibration method [19, 20, 21, 22, 23], we aim to globally calibrate all optical components on the PICs to reduce the accumulation of errors and avoid complex algorithms for decoupling the phase shifters. The most challenging aspect of global calibration is determining the (near-)optimal solution for all parameters that need to be calibrated. To overcome this difficulty, we resort to machine learning to find the optimal parameters. Different from the traditional black-box machine learning approach [42, 43], we utilize a custom-architecture network to build a virtual duplicate of the PICs and then learn the circuit’s real physical process, as illustrated in Fig. 1.
@@
The learning process necessitates a number of training data, comprising randomly generated current distributions $\{\bm{I}\}_{n}$ applied to all the phase shifters and the target probability distributions $\{\bm{P}_{\text{tar}}\}_{n}$ of the output ports on the chip. The index $n=1,\cdots,N$ represents each training sample, and $N$ is the total number of training datasets. We set the current distribution $\bm{I}$ as inputs, which are then fed to the input layer of our custom network. The custom-network-parameterized unitary process $\hat{U}_{\text{ML}}(\bm{a},\bm{b},\bm{\theta})$ can be determined using Eqs. 1 and 2. For a given initial state |$\ket{\psi_{\text{i}}}$, the final state is given by |$\ket{\psi_{\text{f}}}=\hat{U}_{\text{ML}}\ket{\psi_{\text{i}}}$, and the probability distribution $|\langle v|\psi_{\text{f}}\rangle|^{2}$ can be obtained. Considering the relative fiber-coupling efficiency $\bm{\eta}$ between the output ports of the PICs, the final probability distribution is normalized, and then the parameterized network distribution $\bm{P}_{\text{ML}}(\bm{\lambda})$ is generated.
@@
We define the loss function $\mathcal{L}$ as the summation of the L1-norm distance [44] between the target probability distributions in the training datasets $\{\bm{P}{{}_{\text{tar}}}\}_{n}$ and the parameterized network distributions $\{\bm{P_{\text{ML}}(\bm{\lambda})}\}_{n}$: $\mathcal{L}(\bm{\lambda})=\sum_{n}{1\over 2}|\bm{P{{}_{\text{ML}}}}(\bm{\lambda})-\bm{P_{\text{tar}}}|$. The aim of training the custom network is to find the optimal parameters $\bm{\lambda}^{\ast}$ by minimizing the loss function such that the parameterized network distributions approximate the target ones. Then, the parameterized unitary process $\hat{U}_{\text{ML}}$ together with the relative coupling efficiency $\bm{\eta}^{\ast}$ for the trained network can well reflect the target physical process that generates the distributions $\{\bm{P_{\text{ML}}}(\bm{\lambda}^{\ast})\}_{n}\simeq\{\bm{P}{{}_{\text{tar}}}\}_{n}$.
@@ II.2 Model of on-chip quantum walks
$\hat{C}=\sum_{x}\ket{x}\bra{x}\otimes\matrixquantity(\cos\theta&e^{i\alpha}\sin\theta\\ -e^{i\beta}\sin\theta&e^{i(\alpha+\beta)}\cos\theta),$ (3)
@@
where $\cos^{2}\theta$ and $\sin^{2}\theta$ denote the probabilities of a walker moving backwards and forwards on the 1D line, respectively. Parameters $\alpha$ and $\beta$ introduce the relative phase difference. The coin operator $\hat{C}$ can be position- and time-dependent, a feature necessary for specific tasks in QW-based quantum simulations [36] and quantum algorithms [14].
For the QW PIC, the coin state {$\{\ket{\uparrow},\ket{\downarrow}\}$ in the QW model can be cast on the path state (inset of Fig. 2): the upper path of the beamsplitter corresponds to |$\ket{u}$, while the lower path corresponds to |$\ket{l}$. The localized coin state at each position is described by |$\ket{c}=c_{1}\ket{u}+c_{2}\ket{l}$, which can be interpreted as a superposition state located on the two paths of the corresponding beamsplitter. Each beamsplitter with a splitting ratio of 50:50 and phase shifter work together (black circle in Fig. 2) to realize the coin operator $\hat{C}$ in the sense of path states, corresponding to Eq. (3) with $\theta=\pi/4$ and $\beta=0$. The interwoven mesh architecture of the PIC, composed of the on-chip waveguides (black solid lines), realizes the shift operators $\hat{S}$. After $T$-step unitary dynamics of the QW PIC, the size of the position space is $T$, with each position having two paths representing the coin space. Therefore, the total number of output ports of the QW PIC is $2T$.
@@ II.3 Numerical benchmarks
Figure 3: Benchmarking global calibration method using synthetic datasets. (a)$\sim$(d) exhibite the calibration values of $\bm{a}$, $\bm{b}$, $\sin^{2}\bm{\theta}$, and $\bm{\eta}$, respectively, given different numbers of training samples. Black solid lines in (a-d) are the preset target values in theory. The indices of the phase shifters and beamsplitters on the QW PIC are arranged in order from the first column on the far left to the last column, with each column arranged from top to bottom. The indices of the output ports are ordered from top to bottom. The significant deviations in the relative coupling efficiencies $\bm{\eta}$ of the edge output ports at the top and bottom positions are attributed to the extremely low output probability (nearly $10^{-4}$) at these four ports. (e) The values of the loss function versus the number of epochs. (f) Histogram of the values of L1-norm distance for 1000 test samples.
Before training the custom network with noisy experimental results, we begin by numerically benchmarking the calibration method using synthetic datasets. To effectively simulate noisy experimental environments and assess the robustness of our calibration method against external noise, we introduce a type of Gaussian noise into the generated process of the synthetic datasets. Herein, the noise is added to the real-valued amplitudes and phases of the output state |$\ket{\psi_{\text{f}}}$ to ensure the output probability $\bm{P_{\text{th}}}$ remains non-negative. The noise is sampled from a zero-mean Gaussian distribution $\mathcal{N}(0,\epsilon)$ with a standard deviation of $\epsilon=0.01$ defining the noise degree. Then, the calibration process aims to minimize the loss function $\mathcal{L}(\bm{\lambda})=\sum_{n}\frac{1}{2}|\bm{P}{{}_{\text{ML}}}(\bm{\lambda})-\bm{P}{{}_{\text{th}}}|$ for given datasets of noisy synthetic probability distributions $\{\bm{P}_{\text{th}}\}_{n}$.
@@
During the iteration process, simultaneously updating all network parameters using the gradient-descent method is susceptible to getting trapped in a local minimum of $\mathcal{L}$. To mitigate this issue and escape local minima, we utilize a non-simultaneous parameter update method, updating only partial parameters at each epoch while keeping the remaining parameters fixed. The relative coupling efficiency $\bm{\eta}$ is configured to be constantly updated, and the reflectivity parameter $\bm{\theta}$ is frozen at first. We then alternatively train one of the two parameters $\{\bm{a},\bm{b}\}$ while the remaining one is kept frozen and untrainable. Within a single alternating training process, the training for each parameter ceases upon meeting the cut-off condition, where five consecutive decreases in the loss function between adjacent epochs are each less than $10^{-5}$. After eight rounds of alternating training, we proceed to train the last parameter, $\bm{\theta}$, while maintaining the fixed values of $\bm{a}$ and $\bm{b}$. In the end, following the completion of the non-simultaneous parameter updates described above, we proceed to simultaneously update all parameters $\bm{\lambda}$ until the cut-off condition is met.
@@
We examine the impact of varying training sample numbers $N$ on the performance of our calibration method, as depicted in Fig. 3. The optimal parameters $\bm{\lambda}^{\ast}$ of the trained custom network are illustrated in Figs. 3(a)-3(d) for $N=200$ (blue), 300 (orange), 400 (green), and 500 (red). It is observed that as the number of training samples increases, the values of the optimal parameters approach the preset target ones. The corresponding training process of the custom network is displayed in Fig. 3(e), where the loss function is plotted as a function of the training epochs. Compared to the slow-descent regions, each distinctly sharp drop in the loss function occurs as a result of one round of the alternating training of $\bm{a}$ and $\bm{b}$. To quantify the performance of the calibrated PIC, we compute the L1-norm distance $d(\bm{\lambda}^{\ast})={1\over 2}|\bm{P_{\text{ML}}}(\bm{\lambda}^{\ast})-\bm{P_{\text{tar}}}|$ between the target distributions and the network’s predictions for 1000 test samples. The L1-norm distance ranges between 0 for two identical distributions and 1 for complete mismatch. Fig. 3(f) shows the histogram of the L1-norm distances for the test datasets. We can see that nearly all the values of the distances are less than 0.05, indicating that the QW PIC with a circuit depth of 12 can be well calibrated by using several hundred training data samples. Furthermore, our calibration method requires only about a minute for the entire training and validation processes of the network on a laptop equipped with an AMD Ryzen 7 5800H CPU. The training and validation processes are performed using the PyTorch package for Python program [50].

### v2

@@ Abstract
The advancing maturity of photonic integrated circuit (PIC) fabrication technology enables the high integration of an increasing number of optical components onto a single chip. With the incremental circuit complexity, the calibration of active phase shifters in a large-scale PIC becomes a crucially important issue. The traditional one-by-one calibration techniques encounter significant hurdles with the propagation of calibration errors, and achieving the decoupling of all phase shifters for independent calibration is not straightforward. To address this issue, we propose a global calibration approach for large-scale PIC. Our method utilizes a custom network to simultaneously learn the nonlinear phase-current relations for all thermo-optic phase shifters on the PIC by minimizing the negative likelihood of the measurement datasets. Moreover, the reflectivities of all static beamsplitter components can also be synchronizedly extracted using this calibration method. As an example, a quantum walk PIC with a circuit depth of 12 is calibrated, and a programmable discrete-time quantum walk is experimentally demonstrated. These results will greatly benefit the applications of large-scale PICs in photonic quantum information processing.
@@ I Introduction
The traditional calibration method requires the decoupling of all phase shifters followed by individual calibration for each one [19, 20, 21, 22, 23]. However, this approach isn’t ideal for large-scale PICs due to two main challenges. Firstly, while the traditional method can be used to calibrate the standard triangular “Rech” PICs [24, 25], rectangular “Clements” PICs [26, 27, 28, 29], and PICs with cascaded phase shifters [30, 31] by following a specific decoupling order, it is difficult to find an effective and direct way to decouple all phase shifters in the intricate mesh structures. Secondly, deviations in the calibration values of phase shifters in preceding levels may also spread errors in the calibration of subsequent phase shifters. As a result, calibration errors accumulate progressively with the increasing number of phase shifters, eventually becoming significant and non-negligible.
To address these issues, we present an efficient calibration method developed to globally calibrate all optical components on the PICs, thereby mitigating the accumulation of errors inherent in the traditional one-by-one calibration approach. The global calibration method utilizes a custom network with simulation layers that creates a virtual duplicate of the physical PICs, modelling their practical non-unitary processes involving optical loss. Using the gradient-descent approach, we can determine the optimal values of parameters that need to be calibrated, ensuring that the outputs of the network closely resemble the distributions of measurement data. As an example, on a large-scale quantum walk (QW) PIC with a circuit depth of 12, we find that the trained custom network enables the simultaneous extraction of not only the nonlinear phase-current relations of all the thermo-optic phase shifters but also the reflectivities of all the beamsplitters. Moreover, we further experimentally demonstrate the silicon PIC calibration and implement the on-chip QW [32, 33, 34, 35, 36, 37, 38, 39, 40].
This article is organized as follows: In Sec. II, we will introduce our custom-architecture global calibration method, provide a concise overview of the QW PIC, and evaluate the performance of our calibration method using synthetic datasets featuring Gaussian noise. In Sec. III, we will showcase the calibration results of a silicon PIC utilizing noisy experimental datasets and discuss the performance regarding the experimental implementation of QW. Finally, Sec. IV presents the summary of our work and engages in a discussion.
@@ II Theoretical Idea
In this section, we will detail our method for achieving precise and fast calibration of PICs. Our method is based on a custom network with simulation layers situated between the input and output layers, as depicted in Fig. 1. The entire custom network is trained using gradient-based optimization techniques to align its virtual process with the real physical process of the PICs, thereby allowing it to estimate the chip parameters that need to be calibrated.
II.1 Global calibration method
Figure 1: Illustration of the custom architecture for global calibration. The input data consists of the current distribution applied to the thermo-optic phase shifters. The first simulation layer of our custom network transforms the distribution of the current $\bm{I}$ into the distribution of the phase shift $\bm{\phi}$, with a subsequent layer outputting the parameterized final state $\rho_{f}$ for a given initial state $\rho_{i}$. The third simulation layer provides the corresponding measurement probability distribution of the final state $\text{Tr}[\left|v\right\rangle\!\!\left\langle v\right|\rho_{f}]$ in the computational basis |$\ket{v}$. The fourth simulation layer introduces a parameter $\bm{\eta}$ into this probability distribution, which accounts for the relative fiber-chip coupling efficiency between output ports. The last layer renormalizes the probability distribution so that the sum of the probabilities over all output ports equals one, and a parameterized network distribution $\bm{P{{}_{\text{ML}}}}(\bm{\lambda})$ is generated thereafter.
@@ II.1 Global calibration method
where $\sin^{2}\theta$ denotes the reflectivity of the beamsplitter. Considering the inevitable optical loss in a practical chip, the transfer matrix for a lossy BS becomes non-unitary and is given by [41]:
$\hat{U}_{\text{LBS}}=\matrixquantity(\sqrt{\alpha}&0\\ 0&\sqrt{\alpha})\matrixquantity(\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta),$ (2)
where $\alpha$ denotes the transfer efficiency of a single BS, and $1-\alpha$ represents its optical loss. The phase shifter provides a relative phase difference $\phi$ between the two output paths of the beamsplitter, and its matrix representation is $\hat{U}_{\text{PS}}=\text{diag}\{1,e^{i\phi}\}$. For the widely-used thermo-optic phase shifter, the relative phase $\phi$ can be tuned by the current $I$ applied to the phase shifter, according to the following formula [13]:
@@
where $a$ and $b$ are hardware parameters to be determined. Then, the non-unitary process $\hat{U}(\bm{a},\bm{b},\bm{\theta},\bm{\alpha})$ of PICs is the matrix product of its individual components. Moreover, considering the mode-field diameter mismatch between the output ports of the PICs and the optical fiber arrays, it is necessary to introduce another parameter, $\bm{\eta}$, to indicate the fiber-chip coupling efficiency. Consequently, $\bm{\lambda}=\{\bm{a},\bm{b},\bm{\theta},\bm{\alpha},\bm{\eta}\}$ are all undetermined parameters for a linear photonic circuit.
In contrast to the traditional one-by-one localized calibration method [19, 20, 21, 22, 23], we aim to globally calibrate all optical components on the PICs to reduce the accumulation of errors and avoid complex algorithms for decoupling the phase shifters. To achieve this, we utilize a custom network with simulation layers that builds a virtual duplicate of the physical PICs, as illustrated in Fig. 1. The simulation layers of the network leverage prior knowledge of the structure of the PICs to be calibrated. As a result, by utilizing the probability distributions of output ports, global calibration of the PICs could be implemented, thereby mitigating the high demand for complex quantum state and process tomography.
@@
The learning process necessitates a number of training data, comprising randomly generated current distributions $\{\bm{I}\}_{n}$ applied to all the phase shifters and the target probability distributions $\{\bm{P}_{\text{tar}}\}_{n}$ of the output ports on the chip. The index $n=1,\cdots,N$ represents each training sample, and $N$ is the total number of training datasets. We set the current distribution $\bm{I}$ as inputs, which are then fed to the input layer of our custom network. The custom-network-parameterized non-unitary process $\hat{U}_{\text{ML}}(\bm{a},\bm{b},\bm{\theta},\bm{\alpha})$ can be determined using Eqs. (2) and (3). For a given initial state $\rho_{i}$, the final state is given by $\rho_{f}=\hat{U}_{\text{ML}}\rho_{i}\hat{U}_{\text{ML}}^{\dagger}$, and the probability distribution $\text{Tr}[\left|v\right\rangle\!\!\left\langle v\right|\rho_{f}]$ can be obtained. Considering the relative fiber-coupling efficiency $\bm{\eta}$ between the output ports of the PICs, the final probability distribution is normalized, and then the parameterized network distribution $\bm{P}_{\text{ML}}(\bm{\lambda})$ is generated.
@@
We define the cost function $\mathcal{L}$ as the summation of the L1-norm distance [42] between the target probability distributions in the training datasets $\{\bm{P}{{}_{\text{tar}}}\}_{n}$ and the parameterized network distributions $\{\bm{P_{\text{ML}}(\bm{\lambda})}\}_{n}$: $\mathcal{L}(\bm{\lambda})=\sum_{n}\frac{1}{2}|\bm{P{{}_{\text{ML}}}}(\bm{\lambda})-\bm{P_{\text{tar}}}|$. The aim of training the custom network is to find the optimal parameters $\bm{\lambda}^{\ast}$ by minimizing the cost function such that the parameterized network distributions approximate the target ones. Then, the parameterized non-unitary process $\hat{U}_{\text{ML}}$ together with the relative coupling efficiency $\bm{\eta}^{\ast}$ for the trained network can well reflect the target physical process that generates the distributions $\{\bm{P_{\text{ML}}}(\bm{\lambda}^{\ast})\}_{n}\simeq\{\bm{P}{{}_{\text{tar}}}\}_{n}$.
@@ II.2 Model of on-chip quantum walks
$\hat{C}=\sum_{x}\ket{x}\bra{x}\otimes\matrixquantity(\cos\theta&e^{i\gamma}\sin\theta\\ -e^{i\beta}\sin\theta&e^{i(\gamma+\beta)}\cos\theta),$ (4)
$\hat{M}=\sum_{x}\ket{x}\bra{x}\otimes\matrixquantity(\sqrt{\alpha}&0\\ 0&\sqrt{\alpha}),$ (5)
@@
Here, the parameter $\alpha$ determines the loss strength and varies depending on both the position and time step. $\cos^{2}\theta$ and $\sin^{2}\theta$ denote the probabilities of a walker moving backwards and forwards on the 1D line, respectively. Parameters $\beta$ and $\gamma$ introduce the relative phase difference. The coin operator $\hat{C}$ can be position- and time-dependent, a feature necessary for specific tasks in QW-based quantum simulations [35] and quantum algorithms [14].
For the QW PIC, the coin state {$\{\ket{\uparrow},\ket{\downarrow}\}$ in the QW model can be cast on the path state (inset of Fig. 2): the upper path of the beamsplitter corresponds to |$\ket{u}$, while the lower path corresponds to |$\ket{l}$. The localized coin state at each position is described by |$\ket{c}=c_{1}\ket{u}+c_{2}\ket{l}$, which can be interpreted as a superposition state located on the two paths of the corresponding beamsplitter. Each beamsplitter with a splitting ratio of 50:50 and phase shifter work together (black circle in Fig. 2) to realize the coin operator $\hat{C}$ in the sense of path states, corresponding to Eq. (4) with $\theta=\pi/4$ and $\beta=0$. The interwoven mesh architecture of the PIC, composed of the on-chip waveguides (black solid lines), realizes the shift operators $\hat{S}$. After $T$-step non-unitary dynamics of the QW PIC, the size of the position space is $T$, with each position having two paths representing the coin space. Therefore, the total number of output ports of the QW PIC is $2T$.
@@ II.3 Numerical benchmarks
Figure 3: Calibration results using synthetic datasets. (a)$\sim$(e) exhibit the calibration values of $\bm{a}$, $\bm{b}$, $\sin^{2}\bm{\theta}$, $\bm{\alpha}$, and $\bm{\eta}$, respectively, given different numbers of training samples. Black solid lines in (a-e) are the preset target values in theory. The indices of the phase shifters and beamsplitters on the QW PIC are arranged in order from the first column on the far left to the last column, with each column arranged from top to bottom. The indices of the output ports are ordered from top to bottom. The significant deviations in the relative coupling efficiencies $\bm{\eta}$ of the edge output ports at the top and bottom positions are attributed to the extremely low output probability (nearly $10^{-4}$) at these four ports. (f) The values of the cost function versus the number of epochs.
Figure 4: Benchmarking the performance of the global calibration method. Histogram of the values of (a) L1-norm distance, (b) infidelity and (c) purity error for 1000 test samples, given different numbers of training samples.
Before training the custom network with noisy experimental results, we begin by numerically benchmarking the global calibration method using synthetic datasets. To effectively simulate noisy experimental environments and assess the robustness of our calibration method against external noise, we introduce Gaussian noise into the data generation process for the synthetic datasets [48]. Here, the Gaussian noise is added to the output probabilities to mimic experimental measurement errors, such as small vibrations of the optical fibers coupled to the output ports of the PICs. The noise is sampled from a zero-mean Gaussian distribution $\mathcal{N}(0,\epsilon)$ with a standard deviation of $\epsilon=0.01$ defining the noise degree. Then, the calibration process aims to minimize the cost function $\mathcal{L}(\bm{\lambda})=\sum_{n}\frac{1}{2}|\bm{P}{{}_{\text{ML}}}(\bm{\lambda})-\bm{P}{{}_{\text{th}}}|$ for given datasets of noisy synthetic probability distributions $\{\bm{P}_{\text{th}}\}_{n}$.
@@
During the iteration process, simultaneously updating all network parameters using the gradient-descent method is susceptible to getting trapped in a local minimum of $\mathcal{L}$. To mitigate this issue and escape local minima, we utilize a non-simultaneous parameter update method, updating only partial parameters at each epoch while keeping the remaining parameters fixed. The relative coupling efficiency $\bm{\eta}$ is configured to be constantly updated. The reflectivity parameter $\bm{\theta}$ and the relative transfer efficiency $\bm{\alpha}$ are frozen at first. We then alternatively train one of the two parameters $\{\bm{a},\bm{b}\}$ while the remaining one is kept frozen and untrainable. Within a single alternating training process, the training for each parameter ceases upon meeting the cut-off condition, where five consecutive decreases in the loss function between adjacent epochs are each less than $10^{-5}$. After eight rounds of alternating training, we proceed to train the last two parameters, $\bm{\theta}$ and $\bm{\alpha}$, while maintaining the fixed values of $\bm{a}$ and $\bm{b}$. In the end, following the completion of the non-simultaneous parameter updates described above, we proceed to simultaneously update all parameters $\bm{\lambda}$ until the cut-off condition is met.
@@
We examine the impact of varying training sample numbers $N$ on the performance of our calibration method, as depicted in Fig. 3. The optimal parameters $\bm{\lambda}^{\ast}$ of the trained custom network are illustrated in Figs. 3(a)-3(e) for $N=200$ (blue), 300 (orange), 400 (green), and 500 (red). It is observed that as the number of training samples increases, the values of the optimal parameters approach the preset target ones. The corresponding training process of the custom network is displayed in Fig. 3(f), where the cost function is plotted as a function of the training epochs. Compared to the slow-descent regions, each distinctly sharp drop in the cost function occurs as a result of one round of the alternating training of $\bm{a}$ and $\bm{b}$.
To quantify the performance of the calibrated PIC, we first compute the L1-norm distance $d(\bm{\lambda}^{\ast})=\frac{1}{2}|\bm{P_{\text{ML}}}(\bm{\lambda}^{\ast})-\bm{P_{\text{tar}}}|$ between the target distributions and the network’s predictions for 1000 test samples. Furthermore, to fully characterize the output states, we also provide infidelity $F_{\text{in}}=1-\text{Tr}(\sqrt{\sqrt{\rho_{\text{tar}}}\rho_{\text{ML}}(\bm{\lambda}^{\ast})\sqrt{\rho_{\text{tar}}}})$ as a measurement of the distance between the target and network-predicated output quantum states. The values of the L1-norm distance and infidelity range from 0, indicating a perfect match, to 1, indicating a complete mismatch. Figs. 4(a)-4(c) show the histograms of the L1-norm distance, infidelity and purity error $|\text{Tr}[\rho^{2}_{\text{ML}}(\bm{\lambda}^{\ast})]-\text{Tr}[\rho^{2}_{\text{tar}}]|$ for the test datasets, using different numbers of training samples $N$, respectively. We can see that nearly all the values of the three quantities are less than $1.5\times 10^{-2}$, $1\times 10^{-3}$, and $2\times 10^{-15}$ for $N=500$ (red bars), respectively. These results indicate that the QW PIC with a circuit depth of 12 can be well calibrated by using several hundred training data samples. We note that our calibration method requires only about a minute for the entire training and validation processes of the network on a laptop equipped with an AMD Ryzen 7 5800H CPU. The training and validation processes are performed using the PyTorch package for Python program [50].